Solve the systen by the eliination method. (Enter your answers as a comma-separated list. If the syste is inconsistent, enter INCONSISTENT. If the syster is dependent, enter DEPENDENT.) 3x y 19 x2y 13 Show My Work (optional 3. -1 points BerFinMah1 3.1.005 My Notes Solve the system by the elimination method. (Enter your answers as a comma-separated list. If the system is inconsistent, enter INCONSISTENT, If the system is dependent, enter DEPENDENT.) Ask Your Teacher x 4y 22 x, v) D Show My Work (optional
Answer #1

The solution to the given question is provided below for your reference

1.
6x+2y=38
x+2y=13
=>5x=25
=>x=5 and y=4
(x,y)=(5,4)

2.
2x+4y=20
3x+4y=22
=>x=2 and y=4
(x,y)=(2,4)